# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v2

@@ II.2 Weyl semimetal
where the Hamiltonian describes the Weyl semimetal phase, which can be interpreted as a stack of 2D materials with orbital-dependent hopping $t_{z}$, as illustrated in Fig. 1(a2). The matrix form of this Hamiltonian (7) in real space, with OBCs in the $z$-direction and PBCs in the $x$- and $y$-directions, is provided in Subsection SIII A in the Supplemental Material Sup . Similarly, the matrix form with OBCs in the $y$-direction and PBCs in the $x$- and $z$-directions is given in Subsection SIII B in the Supplemental Material Sup .
Rewriting Eq. (7) in the form ${\cal H}_{\rm W}({\bf k})\!=\!\mathbf{h}({\bf k})\cdot{\boldsymbol{\sigma}}$, where $\mathbf{h}({\bf k})\!=\!(h_{x},h_{y},h_{z})$, we define the components as: $h_{x}\!=\!t_{x}\sin(k_{x}a)$, $h_{y}\!=\!t_{y}\sin(k_{y}a)$, $h_{z}\!=\!M_{z}(k_{z})\!-\!m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$, with $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)$. The eigenenergies of this 2D model Hamiltonian (7) are given by
@@
Thus, the energy gap between the top and bottom surfaces at $k_{x}\!=\!k_{y}\!=\!0$ is: $\Delta_{\rm W}\!=\!E_{{\rm W},+}\!-\!E_{{\rm W},-}\!=\!2|M_{z}(k_{z})|$ as shown in Fig. 1(d2).
The Chern number of the Weyl semimetal model (7) can then be calculated using the expression Thouless et al. (1982); Fukui et al. (2005); Xiong et al. (2016); Liu et al. (2019); Lu et al. (2010); Shan et al. (2010); Qin et al. (2022a)
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{B.Z.}}\!\!d^{2}{\bf k}_{||}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\!\cdot\![\partial_{k_{x}}\mathbf{h}({\bf k})\!\times\!\partial_{k_{y}}\mathbf{h}({\bf k})],$ (9)
where ${\bf k}_{||}\!=\!(k_{x},k_{y})$ and the integral is over the 2D Brillouin zone. A detailed analytical expression for the Chern number (9) of the Weyl semimetal model (7) is provided in Section SIV in the Supplemental Material Sup .
For the special case where $k_{x}\!=\!k_{y}\!=\!0$, the eigenenergies of the Weyl semimetal Hamiltonian (7) are:
@@
In this case, the system hosts a pair of gapless Weyl nodes at $(0,0,k_{c\pm})$ with the positions $k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(m_{z}/t_{z})$, under the condition $|m_{z}/t_{z}|\leqslant 1$ as shown in Figs. 1(b2) and 1(c2). These Weyl nodes $k_{c}$ are obtained by solving $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)\!=\!0$ with the condition $-1\leqslant m_{z}/t_{z}\leqslant 1$. For $k_{z}$ values within the range $k_{c-}<k_{z}<k_{c+}$, the system behaves as a Chern insulator with Chern number $C\!=\!1$. Outside this range, for $k_{z}<k_{c-}$ or $k_{z}>k_{c+}$, the system becomes a normal insulator with $C\!=\!0$, as depicted in Fig. 1(c2).
@@ II.3 Comparison between the two models
As shown in Fig. 1, the two models – the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column) – can be viewed as duals of each other. The Weyl-like model in Eq. (2) provides a generic framework for describing Weyl semimetals, including key features such as band touching, Chern number, and Fermi arcs, all of which are captured by the Weyl semimetal model in Eq. (7). Figure 2 illustrates that the two Weyl-like nodes are separated by the bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, with each half of the system hosting a single Weyl-like node. This contrasts with conventional Weyl points, which, due to the no-go theorem Nielsen and Ninomiya (1981a, b), typically appear in pairs in momentum space with opposite chiralities. Below, we will focus on the case involving Floquet engineering.
III Floquet Hamiltonian
@@
III.1 Floquet Weyl-like model
@@
$\frac{z_{c\pm}}{\lambda_{z}}\!=\!\frac{1}{2\pi}[\pi\pm\arccos(\tilde{m}_{z}(A_{0})/V_{z})],$ (18)
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/V_{z}\leqslant 1$. By substituting the model parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $a\!=\!1$ nm into the condition $|\tilde{m}_{z}(A_{0})/V_{z}|<1$, we have $A_{0}<1.034$ nm-1. From Eq. (18), it is clear that the position of the Weyl-like nodes can be tuned by adjusting the intensity of the optical field.
@@
III.2 Floquet Weyl semimetal
@@
$k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(\tilde{m}_{z}(A_{0})/t_{z}),$ (23)
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/t_{z}\leqslant 1$. Substituting typical model parameters, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $a\!=\!1$ nm, into the condition $|\tilde{m}_{z}(A_{0})/t_{z}|<1$, we obtain the constraint $A_{0}<1.034$ nm-1. From Eq. (23), we can conclude that the position of the Weyl nodes can be tuned by adjusting the intensity of the optical field.
@@ III.3 Validity
IV Numerical results for Floquet engineering
@@
IV.1 Floquet Weyl nodes
Figure 3: (a) Floquet Weyl-like nodes and (b) Floquet Weyl nodes as a function of light intensity $A_{0}$. The parameters are $\varphi=0$ (linearly polarized laser), $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
As shown in Fig. 3, the positions of the Floquet Weyl-like nodes [Fig. 3(a)] and the Floquet Weyl nodes [Fig. 3(b)] can be controlled by the light intensity $A_{0}$. Particularly, when $A_{0}$ reaches a threshold value of approximately $\sim 1.034$ nm-1, the two Floquet Weyl-like nodes (or Floquet Weyl nodes) merge into a single node, signaling a topological transition.
IV.2 Floquet Chern number
Figure 4: Chern numbers for (a1,b1) the Floquet Weyl-like model and (a2,b2) the Floquet Weyl semimetal. (a1,a2) $A_{0}\!=\!0.1$ nm-1; (b1,b2) $A_{0}\!=\!0.9$ nm-1. The other parameters are the same as those in Fig. 3.

### v3

@@ II.2 Weyl semimetals
where ${\bf k}\!=\!(k_{x},k_{y},k_{z})$. This Hamiltonian describes a Weyl semimetal phase and can be viewed as a stack of 2D layers with orbital-dependent interlayer hopping $t_{z}$, as illustrated in Fig. 1(a2). The matrix representation of Eq. (10) in real space with open boundary conditions (OBCs) along the $z$-direction and periodic boundary conditions (PBCs) along the $x$- and $y$-directions is provided in Subsection SIII A of the Supplemental MaterialSup . Similarly, the matrix form with OBCs in the $y$-direction and PBCs in the $x$- and $z$-directions is given in Subsection SIII B in the Supplemental Material Sup .
We can rewrite Eq. (10) as ${\cal H}_{\rm W}({\bf k})\!=\!\mathbf{h}({\bf k})\cdot{\boldsymbol{\sigma}}$, where the vector $\mathbf{h}({\bf k})\!=\!(h_{x},h_{y},h_{z})$ has the following components: $h_{x}\!=\!t_{x}\sin(k_{x}a)$, $h_{y}\!=\!t_{y}\sin(k_{y}a)$, $h_{z}\!=\!M_{z}(k_{z})\!-\!m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$, with $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)$. The eigenenergies of the Weyl semimetal model then take the form
@@
The Chern number of this model (10) at fixed $k_{z}$ can be calculated using the expression Thouless et al. (1982); Fukui et al. (2005); Xiong et al. (2016); Liu et al. (2019); Lu et al. (2010); Shan et al. (2010); Qin et al. (2022a)
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{B.Z.}}\!\!d^{2}{\bf k}_{||}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\!\cdot\![\partial_{k_{x}}\mathbf{h}({\bf k})\!\times\!\partial_{k_{y}}\mathbf{h}({\bf k})],$
where the integral is over the 2D Brillouin zone. A detailed analytical expression for the Chern number (II.2) is provided in Section SIV in the Supplemental Material Sup .
For the special case $k_{x}\!=\!k_{y}\!=\!0$, the eigenenergies simplify to
@@
so the corresponding bulk energy gap at $k_{x}\!=\!k_{y}\!=\!0$ is given by $\Delta_{\rm W}\!=\!E_{{\rm W},+}\!-\!E_{{\rm W},-}\!=\!2|M_{z}(k_{z})|$ as shown in Fig. 1(d2). In this case, the system hosts a pair of gapless Weyl nodes at $(0,0,k_{c\pm})$ with the positions $k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(m_{z}/t_{z})$, under the condition $|m_{z}/t_{z}|\leqslant 1$ as shown in Figs. 1(b2) and 1(c2). These Weyl points $k_{c\pm}$ are are determined by solving the condition $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)\!=\!0$. For $k_{z}$ in the range $k_{c-}<k_{z}<k_{c+}$, the system behaves as a Chern insulator with Chern number $|C|\!=\!1$. Outside this range, for $k_{z}<k_{c-}$ or $k_{z}>k_{c+}$, the system becomes a normal insulator with $C\!=\!0$, as depicted in Fig. 1(c2). The corresponding Fermi arc is shown in Fig. 1(e2).
Finally, the total 3D Hall conductivity for the Weyl semimetal model (10) is given by Burkov and Balents (2011)
$\sigma_{H}^{3D}\!=\!\int_{-\pi/a}^{\pi/a}\frac{dk_{z}}{2\pi}\sigma_{H}^{2D}(k_{z})\!=\!\frac{e^{2}}{h}\int_{-\pi/a}^{\pi/a}\frac{dk_{z}}{2\pi}C(k_{z}),$ (14)
where $\sigma_{H}^{2D}(k_{z})\!=\!(e^{2}/h)C(k_{z})$ is the $k_{z}$-resolved 2D Hall conductance and $C(k_{z})$ is defined in Eq. (II.2). Using the Weyl node positions $k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(m_{z}/t_{z})$, we obtain
$\sigma_{H}^{3D}\!=\!\frac{e^{2}}{h}\int_{k_{c-}}^{k_{c+}}\frac{dk_{z}}{2\pi}C(k_{z})\!=\!\frac{e^{2}}{ha\pi}\arccos\left(\!\frac{m_{z}}{t_{z}}\!\right)\!.$ (15)
@@ II.3 Comparison between the two models
As shown in Fig. 1, the two models – the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column) – can be viewed as duals of each other. The Weyl-like model in Eq. (1) provides a generic framework for describing Weyl semimetals, including key features such as band touching, Chern number, and Fermi arc, all of which are captured by the Weyl semimetal model in Eq. (10).
Figure 2 illustrates that the two Weyl-like nodes are spatially separated by a bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, with each half of the system hosting a single Weyl-like node. This contrasts with conventional Weyl points, which, due to the no-go theorem Nielsen and Ninomiya (1981a, b), typically appear in pairs in momentum space with opposite chiralities. In the next section, we turn our attention to the effects of interlayer coupling on the Weyl-like model Hamiltonian (1).
III Interlayer coupling effects
We now incorporate the effects of interlayer coupling into our analysis. The energy spectrum is still obtained by calculating the local density of states Chen et al. (2021); Wan et al. (2011); Yang et al. (2011), while the quantum transport properties are captured by evaluating the layer Hall conductance Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024a); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019) in the presence of interlayer coupling. Additionally, we examine how interlayer coupling influences the total 3D Hall conductivity.
The interlayer hopping term is given by
$\Delta H\!=\!\sum_{j_{z}}\eta_{z}\sigma_{z}\left(c_{j_{z}+1}^{{\dagger}}c_{j_{z}}\!+\!c_{j_{z}}^{{\dagger}}c_{j_{z}+1}\right),$ (16)
where $c_{j_{z}}^{{\dagger}}$ ($c_{j_{z}}$) is the electron creation (annihilation) operator on the $j_{z}$-th layer and $\eta_{z}\sigma_{z}$ represents the nearest-neighbor interlayer hopping matrix. Combining Eq. (1) with the interlayer hopping term, the full Hamiltonian incorporating interlayer coupling becomes
$H_{\rm int}\!=\!\sum_{j_{z}}\left[h_{\text{2D}}({\bf k}_{||})\!+\!V(z)\sigma_{z}\right]c_{j_{z}}^{{\dagger}}c_{j_{z}}\!+\!\Delta H.$ (17)
The matrix representation of the Hamiltonian (17) under OBCs along $z$-direction is provided in Subsection SV A in the Supplemental Material Sup .
III.1 Local density of states
Figure 3: Local density of states (LDOS) [Eq. (18)] for different interlayer hopping strengths $\eta_{z}$. The LDOS is indicated by the red color, while the two green vertical dashed lines mark the positions of the Weyl-like points. (a) $\eta_{z}/t_{0}\!=\!0$; (b) $\eta_{z}/t_{0}\!=\!0.05$; (c) $\eta_{z}/t_{0}\!=\!0.1$; (d) $\eta_{z}/t_{0}\!=\!0.2$. The other parameters are $k_{x}\!=\!k_{y}\!=\!0$, $k_{B}T=0.01$ eV, $V_{z}\!=\!1$ eV, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm.
The local density of states is defined as Zhang et al. (2023)
$\rho(z,E,T)\!=\!\sum_{i}\rho_{i}(z)g(E,E_{i},T),$ (18)
where the spatial density distribution $\rho_{i}(z)$ of the $i$th eigenstates $|\psi_{i}(z)\rangle$ is given by
$\rho_{i}(z)\!=\!\langle\psi_{i}(z)|\psi_{i}(z)\rangle,$ (19)
$g(E,E_{n},T)$ is a Lorentzian broadening function defined as
$g(E,E_{n},T)\!=\!\lim_{k_{B}T\to 0^{+}}\frac{k_{B}T}{\pi[(E\!-\!E_{n})^{2}\!+\!(k_{B}T)^{2}]}.$ (20)
Here, $E$ is the probe energy, $k_{B}$ is the Boltzmann constant, and $T$ is the temperatrue.
Furthermore, we numerically calculate the local density of states for different interlayer hopping strengths, as shown in Fig. 3. In the decoupled case with $\eta_{z}\!=\!0$, Fig. 3(a) shows that the local density of states exhibits clear crossings at the two Weyl-like points. The corresponding probability distributions around these points are symmetric, with equal weight at each Weyl-like point.
For weak interlayer coupling, $\eta_{z}/t_{0}\!=\!0.05$, Fig. 3(b) reveals that the local density of states still exhibits sharp crossings at the same Weyl-like points, similar to the decoupled case in Fig. 3(a). For moderate coupling, $\eta_{z}/t_{0}\!=\!0.1$, Fig. 3(c) shows that the local density of states continues to display distinct crossings at the Weyl-like points, indicating that the features observed in the decoupled case persist. These results suggest that the Weyl-like points are robust against weak to moderate interlayer coupling. However, as the interlayer hopping strength $\eta_{z}$ increases further, Fig. 3(d) shows that the local density of states becomes increasingly dispersive compared to the decoupled scenario. This behavior indicates that strong interlayer coupling disrupts the localized character of the Weyl-like points, ultimately leading to their destruction.
III.2 Layer Hall conductance
Figure 4: Layer Hall conductances for different interlayer hopping strengths $\eta_{z}$. The two green vertical dashed lines indicate the positions of the Weyl-like points. (a) $\eta_{z}/t_{0}\!=\!0$; (b) $\eta_{z}/t_{0}\!=\!0.05$; (c) $\eta_{z}/t_{0}\!=\!0.1$; (d) $\eta_{z}/t_{0}\!=\!0.2$; (e) $\eta_{z}/t_{0}\!=\!0.4$; (f) $\eta_{z}/t_{0}\!=\!0.6$. All other parameters are the same as those used in Fig. 3.
To calculate the layer Hall conductance, we substitute the Berry connection from Eq. (5) into Eqs. (4) and (3). The resulting layer Hall conductances for different interlayer hopping strengths $\eta_{z}$ are shown in Fig. 4.
Comparing the decoupled case $\eta_{z}\!=\!0$ [Fig. 4(a)] with the weakly coupled case $\eta_{z}/t_{0}\!=\!0.05$ [Fig. 4(b)] and the moderately coupled case $\eta_{z}/t_{0}\!=\!0.1$ [Fig. 4(c)], we observe that the layer Hall conductance is only slightly modified near the two Weyl-like points. This result demonstrates that the Weyl-like points remain robust under weak and moderate interlayer couplings.
However, as the interlayer hopping strength $\eta_{z}$ increases further [Figs. 4(d)–4(f)], the layer Hall conductance becomes increasingly smooth and continuous in the vicinity of the Weyl-like points, deviating from the sharp features observed in the decoupled case. Physically, this behavior indicates that strong interlayer coupling destroys the Weyl-like points and significantly alters the bulk topology of the system.
III.3 Total 3D Hall conductivity
Figure 5: Total 3D Hall conductivity [Eq. (8)] as a function of the interlayer hopping strength $\eta_{z}$. (a) Full range: $\eta_{z}/t_{0}\in[0,2]$; (b) Enlarged view of the moderate-coupling regime: $\eta_{z}/t_{0}\in[0,0.2]$. Here, $L_{z}\!=\!\lambda_{z}\!=\!61a$, and the other parameters are the same as those used in Fig. 3.
To investigate the effects of interlayer coupling on the total 3D Hall conductivity, we compute it as a function of the interlayer hopping strength $\eta_{z}$, as shown in Fig. 5(a), based on Eq. (8). The results indicate that for $\eta_{z}/t_{0}>0.05$, the total 3D Hall conductivity decreases monotonically as $\eta_{z}$ increases, eventually approaching zero when $\eta_{z}\sim 2t_{0}$. This behavior indicates that strong interlayer coupling can effectively suppress the total 3D Hall conductivity.
In particular, a closer look at the range $\eta_{z}/t_{0}\in[0,0.2]$ [Fig. 5(b)] reveals an approximate plateau within $\eta_{z}/t_{0}\in[0,0.05]$. This suggests that the total 3D Hall conductivity is robust against weak interlayer coupling.
IV Floquet Hamiltonian
@@
IV.1 Floquet Weyl-like model
@@
$z_{c\pm}(A_{0})\!=\!\frac{\lambda_{z}}{2\pi}\left[\pi\pm\arccos\left(\!\frac{\tilde{m}_{z}(A_{0})}{V_{z}}\!\right)\right],$ (28)
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/V_{z}\leqslant 1$. By substituting the model parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $a\!=\!1$ nm into the condition $|\tilde{m}_{z}(A_{0})/V_{z}|<1$, we have $A_{0}<1.034$ nm-1. From Eq. (28), it is clear that the positions of the Weyl-like nodes can be tuned by adjusting the intensity of the optical field.
Furthermore, substituting Eq. (28) into Eq. (9), we can obtain the total 3D Hall conductivity for the Floquet Weyl-like model (24) as
$\sigma_{H}^{3D}(A_{0})\!\approx\!\!\frac{e^{2}\lambda_{z}}{hL_{z}a\pi}\arccos\left(\frac{\tilde{m}_{z}(A_{0})}{V_{z}}\right)\!.$ (29)
@@
IV.2 Floquet Weyl semimetal
@@
$k_{c\pm}(A_{0})\!=\!\pm\frac{1}{a}\arccos\left(\!\frac{\tilde{m}_{z}(A_{0})}{t_{z}}\!\right),$ (34)
where $\tilde{m}_{z}(A_{0})\!=\!m_{z}\!+\!2m_{0}[1\!-\!{\cal J}_{0}(A_{0}a)]$ and $-1\leqslant\tilde{m}_{z}(A_{0})/t_{z}\leqslant 1$. Substituting typical model parameters, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $a\!=\!1$ nm, into the condition $|\tilde{m}_{z}(A_{0})/t_{z}|<1$, we obtain the constraint $A_{0}<1.034$ nm-1. From Eq. (34), we can conclude that the positions of the Weyl nodes can be tuned by adjusting the intensity of the optical field.
Furthermore, substituting Eq. (34) into Eq. (15), we can obtain the total 3D Hall conductivity for the Floquet Weyl semimetal (30) as
$\sigma_{H}^{3D}(A_{0})\!=\!\frac{e^{2}}{ha\pi}\arccos\left(\!\frac{\tilde{m}_{z}(A_{0})}{t_{z}}\!\right).$ (35)
@@ IV.3 Validity
V Numerical results for Floquet engineering
@@
V.1 Floquet Weyl nodes and total 3D Hall conductivity
Figure 6: (a) Floquet Weyl-like nodes [Eq. (28)] and (b) Floquet Weyl nodes [Eq. (34)] as functions of light intensity $A_{0}$. The parameters are: $\varphi\!=\!0$ (linearly polarized laser), $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{z}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
Based on Eqs. (28) and (34), the positions of the Floquet Weyl-like nodes and Floquet Weyl nodes can be plotted, as shown in Fig. 6. Figure 6(a) and (b) illustrate that the locations of these nodes are tunable via the light intensity $A_{0}$. Notably, when $A_{0}$ reaches a critical value of approximately $1.034$ (nm)-1, the two Floquet Weyl-like nodes (or Floquet Weyl nodes) merge into a single point, indicating a topological phase transition.
Figure 7: Total 3D Hall conductivities calculated from Eqs. (29) and (35) for (a) the Floquet Weyl-like model and (b) the Floquet Weyl semimetal, as functions of light intensity $A_{0}$. All other parameters are the same as those in Fig. 6.
In parallel, using Eqs. (29) and (35), the total 3D Hall conductivity as a function of $A_{0}$ is calculated for both the Floquet Weyl-like model and the Floquet Weyl semimetal, as shown in Fig. 7. The results show that the total 3D Hall conductivity decreases monotonically with increasing $A_{0}$, ultimately approaching zero as $A_{0}$ nears the critical value of $\sim 1.034$ (nm)-1.
V.2 Floquet Chern number
Figure 8: Chern numbers for (a1, b1) the Floquet Weyl-like model and (a2, b2) the Floquet Weyl semimetal. Panels (a1, a2) correspond to $A_{0}\!=\!0.1$ (nm)-1, and panels (b1, b2) correspond to $A_{0}\!=\!0.9$ (nm)-1. The two green vertical dashed lines indicate the positions of the Weyl-like or Weyl points. All other parameters are the same as those in Fig. 6.